OpenStudy (anonymous):

one number is five times as large as another number. the difference between them is 120. what is the smaller number?

OpenStudy (anonymous):

need help with these problems too! (image ^)

OpenStudy (zzr0ck3r):

let a be the first number and b the second a=5b that is for one number is 5 times larger than the other a-b=120, this is the difference between them being 120 (assuming a is the bigger one) so a=5b, a-b=120 plug the first equation into the second a-b=120 becomes 5b-b=120 implies 4b=120 implies b=30, which implies a=150

OpenStudy (zzr0ck3r):

so you pick the next one and lets try it